4y + x = 3
2x + 8y = 5

Answer by oberobic(2304)

4y + x = 3
Rearrange
x + 4y = 3
2x + 8y = 5
Multiply first equation by 2
2x + 8y = 6
2x + 8y = 5
Well, that's impossible.
2x + 8y cannot equal 6 and equal 5.
Done.